A 200m long train starts from rest at t=0 with constant acceleration 4cm/s . The head light of the engine is switched ON at t=60s and its tail light is switched ON at t=120s . Find the distance between these two events for an observer standing on platform.
 

Grade:10

2 Answers

Vikas TU
6 years ago
u= 0
s = ut +1/2at2
s = ½ x4x60x60
    = 72 m
s = ½ x4 x120x 120
   = 288m
Difference between two events = 288-72
                                                 = 216 m
Shivanandan D N
4 years ago
Considering the head light to be on the origin, the tail light will be on -200 in the coordinate axis. Initial velocity is equal to 0 m/s. We Know, s = ut + ½ at.
Distance Travelled head light = s1 = (½ x 4 x 60 x 60)/(10 x 10) = 72 m
Distance Travelled tail light    = s2 = (½ x 40 x 60 x 60)/(10 x 10) – 200  = 288 m [ If we consider the tail light to be on the origin we get 288, but we considered it to be on -200. Therefore subtracting 200 from 288].
The distance between two events = s- s1 = 16 m.